The resultant of two forces 3P and 2P is R if the first force is doubled then the resultant is also doubled. the angle between the two forces is 

A

60∘

B

120∘

C

70∘

D

180∘




1
Expert's answer
2021-09-06T00:49:51-0400

"Solve \\space two \\space equation :\\\\\nR ^\n2\n =9P \n^2\n +4P^ \n2\n +12P^ \n2\n cos\u03b8=13P^ \n2\n +12P^ \n2\n cos\u03b8 .....(1)\\\\\n\n (2R)^ \n2\n =(2\u00d73P) ^2\n2\n +4P^ \n2\n +24P^ \n2\n cos\u03b8=40P \n^2\n +24P^ \n2\n cos\u03b8 ....(2)\\\\\nMultiplying (1) \\space by \\space 4, 4R^ \n2\n =52P^\n2\n +48P^2 \n2\n cos\u03b8 .....(3)\\\\\nEquating \\space (2) \\space and \\space (3)\\\\\n40P ^\n2\n +24P^ \n2\n cos\u03b8=52P ^\n2\n +48P^ \n2\n cos\u03b8\\\\\n\u21d212P^ \n2\n =\u221224P^ \n2\n cos\u03b8\\\\\n\u21d2cos\u03b8=\u2212 \n2\n1\n\u200b\n =cos120^ \n\u2218\\\\\n \n\u21d2\u03b8=120^\n\u2218"